# Queryable Prototype Multiple Instance Learning with Vision-Language Models for Incremental Whole Slide Image Classification

### v2

@@ Introduction
As shown in Figure 2, class text feature enhancing branch is designed to generate an enhanced class feature $\boldsymbol{f^{\prime}}_{\text{txt}}$.
@@
IL Type Method ACC ($\uparrow$) Upper-bound Ratio ($\uparrow$) Forgetting ($\downarrow$) BWT ($\uparrow$) Params (M) Masked ACC ($\uparrow$)
@@
Table 1: Main results in class-incremental learning scenario on forward-order training. The best performances are highlighted as bold. “Params” represents the number of learnable parameters. Gray numbers are directly cited from published paper. “/30” and “/100” mean buffer $30$ WSIs and buffer $100$ WSIs, respectively. We present the Masked ACC reflecting task-incremental learning scenario for reference. /∗∗/∗∗∗ denote there are significant differences (paired t-test p-value $<$ $0.05$/$0.01$/$0.001$) between the best performance and others.
@@
(2) Two regularization-based methods are almost ineffective in WSI class-incremental learning scenario. For example, as a regularization-based method, LwF only achieves an ACC of $0.378$$\pm$ $0.081$, which is slightly higher than the lower bound of $0.308$$\pm$ $0.045$, resulting in a very low Upper-bound Ratio of $0.412$$\pm$ $0.083$. Meawhile, EWC also shows almost ineffective in incremental classification.
@@
Tunable Vector Class Ensemble Class Similarity Loss
@@
Table 3: Ablation studies on main components.
Ablation ACC ($\uparrow$) Upper-bound Ratio ($\uparrow$) Forgetting ($\downarrow$) BWT ($\uparrow$)
w/o key 0.675 $\pm$ 0.029 0.742 $\pm$ 0.045 0.332 $\pm$ 0.034 -0.332 $\pm$ 0.034
w/o matching penalty 0.659 $\pm$ 0.029 0.724 $\pm$ 0.037 0.356 $\pm$ 0.048 -0.356 $\pm$ 0.048
QPMIL-VL 0.890 $\pm$ 0.021 0.982 $\pm$ 0.032 0.027 $\pm$ 0.014 -0.027 $\pm$ 0.014
@@
(1) Our QPMIL-VL method is consistently superior to all other methods in both class-incremental and task-incremental learning scenarios, maintaining a SOTA position. (2) Regularization-based methods are still ineffective in class-incremental learning scenario due to suffering from severe catastrophic forgetting. (3) Most rehearsal-based and pre-trained VLM-based methods often achieve good classification performance. Except for MI-Zero and our QPMIL-VL, the best ACC among these two types of methods reaches up to $0.703$$\pm$ $0.049$ with an Upper-bound Ratio of $0.777$$\pm$ $0.053$, achieved by ER-ACE.
Comparison between Forward-order and Reverse-order Training. Additionally, we observe that on reverse-order training, the performance of most methods is noticeably lower than on forward-order training. For example, our QPMIL-VL decreases to an ACC of $0.859$$\pm$ $0.032$. In contrast, its ACC is $0.890$$\pm$ $0.021$ on forward-order training (as shown in Table 1). One possible explanation to this decline is that in incremental learning, later datasets can affect earlier ones. ESCA, with only $150$ slides, is the first to be learned, while NSCLC, with $965$ slides, is the last. NSCLC is dominant in the total number of training samples, which could cause a greater influence on the earlier datasets, leading to more forgetting.
@@
I) on Main Components To analyze the impact of the main components on performance, we conduct a set of ablation studies on forward-order training. The results are shown in Table 3, where the first row represents our baseline (FineTune). By comparison, two obvious findings can be observed as follows.
@@
II) on Prototype Pool Furthermore, we conduct another set of ablation studies specifically on the queryable prototype pool. Our experimental results are shown in Table 4.
@@ A. Dataset Details
The benchmark consists of four public WSI datasets from TCGA repository: non-small cell lung carcinoma (NSCLC), invasive breast carcinoma (BRCA), renal cell carcinoma (RCC) and esophageal carcinoma (ESCA), as shown in Table 5.
@@ B. Metric Details
After completing the incremental training, we can obtain an Accuracy performance matrix, as illustrated in Table 6. Subsequently, the relevant metrics reported in our experiments are calculated using this performance matrix.
@@ C. More Details of Class Ensemble
Table 7 provides all the class names and templates used in our implementation of the class ensemble. They are adapted from those used in CONCH.
@@ E. Query Vector Generation Strategies
MaxPooling vs MeanPooling. To investigate the impact of different pooling operations on the performance of QPMIL-VL when generating query vector, we conduct comparative experiments using MaxPooling and MeanPooling. Table 10 presents the experimental results.

### v3

@@ Introduction
As shown in Fig. 2, class text feature enhancing branch is designed to generate an enhanced class feature $\boldsymbol{f^{\prime}}_{\text{txt}}$.
@@
IL Type Method ACC ($\uparrow$) Upper-bound Ratio ($\uparrow$) Forgetting ($\downarrow$) BWT ($\uparrow$) Params Masked ACC ($\uparrow$)
@@
Table 1: Main results in class-incremental learning scenario on forward-order training. The best performances are highlighted as bold. “Params” represents the number of learnable parameters (M). Gray numbers are directly cited from published paper. “/30” and “/100” mean buffer $30$ WSIs and buffer $100$ WSIs, respectively. We present the Masked ACC reflecting task-incremental learning scenario for reference. /∗∗/∗∗∗ denote there are significant differences (paired t-test p-value $<$ $0.05$/$0.01$/$0.001$) between the best performance and others.
@@
(2) Two regularization-based methods are almost ineffective in WSI class-incremental learning scenario. For example, as a regularization-based method, LwF only achieves an ACC of $0.378$$\pm$ $0.081$, which is slightly higher than the lower bound of $0.308$$\pm$ $0.045$.
@@
TV CE CSL
@@
Table 3: Ablation studies on main components (TV: Tunable Vector, CE: Class Ensemble, CSL: Class Similarity Loss).
Ablation ACC ($\uparrow$) Forgetting ($\downarrow$)
w/o key 0.675 $\pm$ 0.029 0.332 $\pm$ 0.034
w/o matching penalty 0.659 $\pm$ 0.029 0.356 $\pm$ 0.048
QPMIL-VL 0.890 $\pm$ 0.021 0.027 $\pm$ 0.014
@@
(1) Our QPMIL-VL method is consistently superior to all other methods in both class-incremental and task-incremental learning scenarios, maintaining a SOTA position. (2) Regularization-based methods are still ineffective in class-incremental learning scenario due to suffering from severe catastrophic forgetting. (3) Most rehearsal-based and pre-trained VLM-based methods often achieve good classification performance.
Comparison between Forward-order and Reverse-order Training. Additionally, we observe that on reverse-order training, the performance of most methods is noticeably lower than on forward-order training. For example, our QPMIL-VL decreases to an ACC of $0.859$$\pm$ $0.032$. In contrast, its ACC is $0.890$$\pm$ $0.021$ on forward-order training (as shown in Tab. 1). One possible explanation to this decline is that in IL, later datasets can affect earlier ones. ESCA, with only $150$ slides, is the first to be learned, while NSCLC, with $965$ slides, is the last. NSCLC is dominant in the total number of training samples, which could cause a greater influence on the earlier datasets, leading to more forgetting.
@@
I) on Main Components To analyze the impact of the main components on performance, we conduct a set of ablation studies on forward-order training. The results are shown in Tab. 3, where the first row represents our baseline (FineTune). Two obvious findings can be observed as follows.
@@
II) on Prototype Pool Furthermore, we conduct another set of ablation studies specifically on the queryable prototype pool. Our experimental results are shown in Tab. 4.
@@ A. Dataset Details
The benchmark consists of four public WSI datasets from TCGA repository: non-small cell lung carcinoma (NSCLC), invasive breast carcinoma (BRCA), renal cell carcinoma (RCC) and esophageal carcinoma (ESCA), as shown in Tab. 5.
@@ B. Metric Details
After completing the incremental training, we can obtain an Accuracy performance matrix, as illustrated in Tab. 6. Subsequently, the relevant metrics reported in our experiments are calculated using this performance matrix.
@@ C. More Details of Class Ensemble
Tab. 7 provides all the class names and templates used in our implementation of the class ensemble. They are adapted from those used in CONCH.
@@ E. Query Vector Generation Strategies
MaxPooling vs MeanPooling. To investigate the impact of different pooling operations on the performance of QPMIL-VL when generating query vector, we conduct comparative experiments using MaxPooling and MeanPooling. Tab. 10 presents the experimental results.